OpenStudy (anonymous):

if three sides of trapezoid are each 6 inches long, how long must the fourth side be if the area is maximum? ans. is 12

OpenStudy (anonymous):

use the formula for the area of a trapezoid...then to maximize you take the first derivative set it equal to 0, and find critical values...does this sound familiar

OpenStudy (dumbcow):

what class is this for?

OpenStudy (dumbcow):



OpenStudy (dumbcow):

\[A = \frac{1}{2}h(x+6)\] from the right triangle we can define "h" in terms of "x" \[h^{2} +\frac{(x-6)^{2}}{4} = 36\] \[h = \frac{\sqrt{144 - (x-6)^{2}}}{2}\] sub into Area equation \[A = \frac{1}{4}(x+6) \sqrt{144-(x-6)^{2}}\] differentiate and set equal to 0 \[\frac{dA}{dx} = \frac{1}{4}\sqrt{144-(x-6)^{2}} -\frac{x-6}{\sqrt{144-(x-6)^{2}}}=0\] solve for x \[144 - (x^{2}-12x+36) = 4x-24\] \[x^{2} -8x-132 = 0\] \[x = 4+2\sqrt{37}\]

OpenStudy (dumbcow):

sorry i made an error with the derivative...obviously the answer must be 12